OpenStudy (anonymous):

Solve the equation Quantity x plus 1 all over 3 = 2 for x. Please show your work and solution in the space provided.

OpenStudy (anonymous):

\[\frac{ x+1 }{ 3 } =2 \] Multiply \[\frac{ 3 }{ 1 }\] to both sides. Leaving you with x+1 =6. Subtract 1 from both sides, x=5.